OpenStudy (anonymous):

The inequality 2x - 3y ≥ 5 is satisfied by point (, ). True False

OpenStudy (anonymous):

\[\left(\begin{matrix}\left[\begin{matrix}1 & 1 \\ 2 & 3\end{matrix}\right] \\ \end{matrix}\right)\]

OpenStudy (anonymous):

without the brackets

OpenStudy (saifoo.khan):

What you wrote the matrix?

OpenStudy (anonymous):

@saifoo.khan idk its an example i have to do